OpenStudy (anonymous):

'K' stands for a whole number. 'K' is greater than 100. 'K' - 7 is less than 90. List all the numbers that 'K' could be. (I'm stuck so i need all your help..... thnk u danii xxx)

OpenStudy (anonymous):

consider 101. 101-7 = 94 which is greater than 90. So there cant be any number which satisfies the conditions.

OpenStudy (anonymous):

but i thought that the number could be 102 ???

OpenStudy (anonymous):

i meant i thought it could be a number like 95 and soo on because 95-7= 87 and 95+7= 102

OpenStudy (anonymous):

But the question says 'k' should be greater than 100 r8?

OpenStudy (anonymous):

nooo it says k +++++ 7 is greater than 100

OpenStudy (anonymous):

i am not sure if i know what '+++++' means. does it mean k+5*7?

OpenStudy (anonymous):

soz i mean '+'

OpenStudy (anonymous):

can i ask u a diff question plzz

OpenStudy (anonymous):

No need to ask. this is a study group. Shoot.

OpenStudy (anonymous):

M stands for a whole number greater than 10 and less than 20 N stands for a whole number greater than 2 and less than 10 What is the smallest number that M x N could be?

OpenStudy (anonymous):

I am not so sure about this. I feel there is more to this problem. Anyways, Smallest possible value for M is 11 and smallest possible value for N is 3. So, smallest M x N = 33.

OpenStudy (anonymous):

awww thnxxx

OpenStudy (anonymous):

what about the largest number that M-N could be?

OpenStudy (anonymous):

Just do (largest possible M)-(largest possible N)=19-9=10

OpenStudy (anonymous):

so the answer is 10??